A triangle has an angle that measures 85°. The other two angles are in a ratio of 6:13. What are the measures of those two angles?
Answer by ankor@dixie-net.com(22740)

triangle has an angle that measures 85°.
The other two angles are in a ratio of 6:13. What are the measures of those two angles?
:
let x = the multiplier
then
6x = the measure of one angle
13x = measure of the other angle
180 - 6x - 13x = 85
-19x = 85 - 180
-19x = -95
x = -95/-19
x = +5
then
6(5) = 30 degrees is one angle
and
13(5) = 65 degrees is the other
:
;
Check 30 + 65 + 85 = 180